OpenStudy (anonymous):

I really need help with multiplying rational expressions. I just can't seem to understand it. Here is one I need help with: 6a/-1 * 9/-8a^2

OpenStudy (anonymous):

\[\frac{ 6a }{ -1} * \frac{ 9 }{ -8a^2 }\] (easier to read)

OpenStudy (nikato):

There are two ways to approach it 1. Just multiply the numerator and then denominator. Then simplify 6a x 9. 54a -------- = -------- -1 x -8a^2. 8a^2

OpenStudy (nikato):

We know both 54a and 8a^2 are divisible by 2a So you can simplify to 27/ 4a

OpenStudy (nikato):

For the second method, you would reduce and then multiply which is helpful for more complicated stuff Look how 6a and 8a^2 are both divisible by 2a So 3 9. 27 --- x --- = ---- -1. -4a. 4a Same answer but less step/work

OpenStudy (anonymous):

Okay @nikato, that makes sense. but what about a problem like this: \[\frac{ -7x }{ 5y } * \frac{ 5x }{ -7 }\]

OpenStudy (nikato):

Same thing. You can use either but I prefer the second becuz it's easier. Hint: look at diagonal numbers and see what you can simplify In this problem you have -7x and -7 which are both divisible by -7 And you have 5x and 5y which are both divisible by 5 So x. x. x^2 -- * --- = ------- y. 1. y